# GastCoCo: Graph Storage and Coroutine-Based Prefetch Co-Design for Dynamic Graph Processing

### v3

@@ 6.1. Execution Strategy Tuner
In this subsection, we explain how the execution strategy tuner tailors partition and scheduling strategies based on graph processing tasks and how the task allocator allocates tasks to coroutines. It is noteworthy that modeling each task as a coroutine and creating the interleaved execution mode among multiple tasks is feasible and easy, but we focus on building interleaved execution patterns among multiple coroutines within a single task in this subsection. We present the logic of the execution strategy tuner in Figure 8 for a better understanding.
@@ 4 Function Scheduler($coroutine\_pool$):
9                         $coroutine\_pool[i].$destory()
@@ 15 Function Coroutine_Func():
Graph Update Tasks. In GastCoCo, we support both individual updates and batch updates, depending on the specific requirements of the application scenario. However, it is important to note that the coroutine optimization scheme used in GastCoCo can only optimize batch edge updates. Specifically, for individual updates, we apply vertex-level locks on the CBList to ensure accuracy. For batch updates, similar to systems that support batch updates (e.g., Aspen (Dhulipala et al., 2019) and Terrace (Pandey et al., 2021)), to avoid the overhead of locks caused by data conflicts, we classify update tasks by source vertex before updating. Each vertex that is waiting to be updated will have a collection of update tasks. The task allocator models each vertex’s task collection as a coroutine, which is similar to FindNeighbor($edge$) in algorithm 2. Since insertions may modify the structure making it difficult to estimate the number of suspension points before the update, we opt for the trimmed-polling scheduler for batch updates.
@@ 6.2. Hybrid Prefetching
Hardware prefetching fetches data silently to users. However, if hardware prefetching succeeds, repeatedly operating software prefe- tching will lead to additional overhead. Accordingly, the ideal scenario is to use software prefetching when the hardware prefetching fails to fetch data and to avoid using software prefetching when hardware prefetching is effective.
Assume that the cost of a cache miss is $C_{m}$ and the probability of a cache miss being resolved by hardware prefetching is $P_{h}$. This probability varies based on the memory layout of the data. The software prefetching solution via coroutines incurs a certain overhead, $C_{coro}$. If the cost $C_{m}\times(1-P_{h})\textless C_{coro}$, then the software prefetching becomes redundant. To avoid this overhead of redundant data fetching, we aim to avoid employing software prefetching on data that is highly likely to be prefetched by the hardware prefetchers.
We consider the phenomenon caused by the skewed graph structures from three perspectives: ① Array-like structures offer good performance for computation through hardware prefetching, but the frequent update requirements can limit their use. ② Hardware prefetching can also be effective on linked list-like structures, but fetching the front few blocks of a linked list will cause cache misses in theory. When it comes to the graphs where most vertices are low-degree (i.e., the linked lists are short), hardware prefetching has no advantage on most linked lists, thus dramatically increasing the time cost. ③ Maintaining ordered data is also necessary for graph queries and GPM, so tree-like structures are needed. However, hardware prefetching does not work effectively on them.
Therefore, , we aim to use software prefetching to complement hardware prefetching for better performance. However, implementing software prefetching via coroutines has its own overhead, as discussed in section 5. To better utilize both hardware and software prefetching, we design four prefetching strategies, as shown in Figure 9.
@@
Hybrid strategy I (Hybrid prefetching based on the block size). In the implementation of CBList, all small chunks are arranged contiguously in memory, resulting in a high probability of being prefetched by the hardware prefetcher. In this case, we skip software prefetching for small chunks and instead rely on the hardware prefetcher’s inherent capability to prefetch these small chunks. In most cases, this strategy is more effective than the ‘All Soft’ strategy. It is more suitable for scenarios where small chunks have a high cache hit rate.
Hybrid strategy II (Hybrid prefetching based on the hotness). To further prevent redundant data fetching, we consider the cold start issue of the jump-pointer mechanism described in Section 2.2. So based on Hybrid I, we use software prefetching at the beginning of the linked list. For the latter part of the linked list, we only rely on the hardware prefetcher’s jump-pointer mechanism. This strategy utilizes software or hardware prefetching based on the cache’s ”cold” and ”hot” data status. Under ideal conditions, this strategy achieves optimal results. Determining the optimal threshold parameters for using hardware or software prefetching based on the cache’s state requires iterative tuning and adjustments. So it is particularly suitable for scenarios where repeated fine-tuning of computations is possible.
@@ 7. EVALUATION
This section first introduces the experimental setup, and then validates the efficiency of the proposed system through various evaluations.
@@ 7.1. Experimental Setup
We run the experiments on a server equipped with one Intel Xeon Platinum 8269CY CPU whose clock speed is 2.5GHz. The server contains 26 cores (52 hyper-threads) and 371GB of main memory. The sizes of the L1-cache, L2-cache, and L3-cache of the CPUs are 832KB(data cache)+832KB(instruction cache), 26MB, and 35.8MB, respectively. We compiled all systems with the O3 optimization flag using GCC v10.3. All experiments are conducted with 16 worker threads by default unless otherwise stated, with each test running five times to report the average outcome.
Datasets. We used five real-life graphs in our evaluation (see Table Table 2), including social networks Livejournal (Backstrom et al., 2006), Hollywood (Boldi and Vigna, 2004) and Com-friendster (Yang and Leskovec, 2012), and web graph UK-2002 (Boldi and Vigna, 2004) and Orkut (Yang and Leskovec, 2012). The detailed information of the datasets used is shown in Table 2, which includes the name of the dataset, the number of vertices, the number of edges, and the average degrees. To simulate the realism of dynamic graph processing scenarios in practice, the datasets are shuffled. This is because other structures containing pointers may be penalized during loading due to the inherent order of the data.To ensure fairness, all systems load datasets using the weighted graph mode. Even if the dataset is an unweighted graph, random weights are generated. Additionally, we used the SNB-sf1000 dataset (Angles et al., 2020) to test the system’s update performance in real-world scenarios. The SNB-sf1000 dataset includes timestamped person nodes and their relationships, along with 7,285 vertex deletions and 39,877,751 edge insertions. Ogbl-citation2 (Wang et al., 2020) dataset, which includes 128-dimensional word2vec features, was used to test graph attribute updates.
Workloads. We conduct three types of workloads. (i) Graph Query: we randomly query 5% edges of the data set. (ii) Graph Algorithms: we use five typical graph analysis algorithms, i.e., BFS, Single Source Shortest Path (SSSP) (Chakaravarthy et al., 2017), PageRank (PR) (Page et al., 1998), Connected Components (CC) and Label Propagation (LP). (iii) Graph Update: we synthetically generate 10,000 of random graph updates in form of edge insertions and edge deletions.
@@
LLAMA, Teseo and PCSR utilize structures similar to CSR to enhance data locality. LiveGraph employs a log-structured approach for data insertion to achieve faster insertion performance. GraphOne and RisGraph adopt block-based adjacency lists and adjacency arrays, respectively, to support rapid data insertion. Sortledton and Terrace employ different structures for storing neighbor data based on vertex degree, striking a balance between insertion performance and graph analysis capabilities.
@@ 7.2. Performance of Graph Query
We first evaluate the performance of edge queries by GastCoCo on all datasets in Table 2, which plays a crucial role in multiple domains such as social network analysis, recommendation systems, and transportation networks. For example, in social networks, edge queries can be utilized to check whether there is a direct friendship between two users.
Figure 10 shows the normalized time for each system to execute queries on different datasets. Here the response time of GastCoCo is treated as the baseline, i.e., finishes in unit time. We can see that GastCoCo outperforms others in all the cases. Specifically, GastCoCo achieves an average 13.7$\times$ (up to 16.8$\times$) speedup over Terrace, 7.4$\times$ (up to 12.0$\times$) speedup over LLAMA, 2.8$\times$ (up to 7.2$\times$) speedup over PCSR, 10.9$\times$ (up to 20.4$\times$) speedup over GraphOne, 4.7$\times$ (up to 8.5$\times$) speedup over RisGraph, 6.0$\times$ (up to 9.1$\times$) speedup over LiveGraph, 4.6$\times$ (up to 6.2$\times$) speedup over Teseo, 5.6$\times$ (up to 7.6$\times$) speedup over Sortledton. For Terrace, LLAMA, and GraphOne, they experienced a slowdown exceeding 10$\times$ on the Hollywood dataset. This can be attributed to the dataset’s vertices having the highest average degree(i.e., 101.81), which leads to significant query overhead for vertices with high degrees due to the storage structures of these three systems during retrieval. Specifically, Terrace adopts a hierarchical structure. When the degree of a vertex is large, its neighbors will be distributed in multiple layers of data structures, which results in the need to search across multiple storage structures when querying. For GraphOne, it uses a block-based adjacency list to store the neighbors of each vertex. Therefore, for vertices with a high degree, this necessitates traversing a longer chain, thereby extending the query time. Compared with other systems, GastCoCo also achieves better performance, mainly because GastCoCo adopts more ”stubby” structures, B+ Tree, as part of our edge storage. The B+ tree, due to its strict rebalance rules, has fewer levels, meaning we encounter fewer pointers and experience fewer cache misses during queries. We can offset the benefits with software prefetching via coroutines to amortize the rebalance overhead during updates.
@@ 7.3. Performance of Graph Analysis
We next evaluate the performance of graph analysis, including the execution time of five graph analytics algorithms, by comparing GastCoCo with competitors. Figure 11 shows the normalized time for each system to execute graph analysis algorithms on different datasets. Here the response time of GastCoCo is treated as the baseline, i.e., finishes in unit time. We can see that GastCoCo outperforms others in most cases. Specifically, GastCoCo achieves an average 1.4$\times$ (up to 3.9$\times$) speedup over Terrace, 41.1$\times$ (up to 243.8$\times$) speedup over LLAMA, 11.0$\times$ (up to 39.7$\times$) speedup over PCSR, 21.7$\times$ (up to 79.4$\times$) speedup over GraphOne, 4.9$\times$ (up to 22.3$\times$) speedup over RisGraph, 4.2$\times$ (up to 16.0$\times$) speedup over LiveGraph, 5.8$\times$ (up to 21.7$\times$) speedup over Teseo, 2.8$\times$ (up to 6.8$\times$) speedup over Sortledton. As shown in Figure 11, LLAMA and GraphOne exhibit the most significant slowdown. LLAMA employs a multi-version array to store the neighboring edges of vertices. As data is inserted, the edges of a vertex may be distributed across multiple arrays, which is disadvantageous for graph analysis algorithms since they always require access to all neighbors of a vertex. LLAMA connects the edges of a vertex distributed in different arrays by pointers, thereby reducing search time. However, this still results in extensive pointer chasing. GraphOne faces similar issues with expensive pointer chasing due to its use of block-based adjacency lists. Furthermore, Figure 11 shows that Terrace, PCSR, and Sortledton achieved better graph analysis performance compared to other competitors. This is attributed to Terrace and Sortledton using arrays to store the neighbors of vertices with low degrees, and PCSR using arrays to store the neighbors of all vertices, leading to better spatial locality. However, this design also results in higher update costs due to data movements caused by insertions/deletions, which will be detailed in Section 7.4. Additionally, Terrace and Sortledton employ B-Trees or skip lists to store the neighbors of high-degree vertices, still leading to extensive pointer chasing. In contrast, GastCoCo utilizes software prefetching and coroutines to accelerate data retrieval in graph analysis and reduce cache misses, thereby achieving superior graph analysis performance.
@@ 7.4. Performance of Graph Updates
We evaluate the graph update performance of different systems on LiveJournal, Hollywood and UK-2002 for continuous edge insertions/deletions of 10K edges. Figure 12 shows the throughput (edges per second) for all systems under different data sets. GastCoCo outperforms all competitors in all tests. Specifically, GastCoCo achieves an average 2.5$\times$ (up to 4.5$\times$) speedup over Terrace, 152.4$\times$ (up to 180.2$\times$) speedup over LLAMA, 33.3$\times$ (up to 35.4$\times$) speedup over PCSR, 4.3$\times$ (up to 5.0$\times$) speedup over GraphOne, 58.2$\times$ (up to 75.0$\times$) speedup over RisGraph, 74.4$\times$ (up to 109.4$\times$) speedup over LiveGraph, 68.8$\times$ (up to 87.2$\times$) speedup over Teseo, 33.3$\times$ (up to 35.4$\times$) speedup over Sortledton. It can be observed that Terrace, GraphOne, and GastCoCo have higher throughput. For GraphOne, it utilizes a block-based adjacency list storage structure, which is particularly advantageous for insertions. However, as demonstrated by previous experiments, it performs poorly in graph queries and analysis. Both Terrace and GastCoCo adopt a batch updates design, i.e., classify update tasks by source vertices before updating, to avoid lock overhead caused by data conflicts. Moreover, since Terrace maintains the updated data in an ordered manner, a search operation is required to determine the position for the new data, leading to a significant number of cache misses. In contrast, although GastCoCo also maintains ordered updates, our coroutine-related design effectively reduces cache misses, thereby achieving better performance than Terrace.
@@ 7.5. Varying Batch Size of Graph Updates
We also conduct experiments to evaluate the impact of batch size of graph updates on throughput performance. Figure 13 reports the throughput of the systems on graph updates with different batch sizes on LiveJournal, Hollywood, Orkut and UK-2002. For each batch size from $10$ to $10^{7}$, we report the throughput (edges per second) for all systems. Figure 13 shows that GastCoCo outperforms all systems with batch sizes ranging from $10^{3}$ to $10^{7}$, primarily for reasons consistent with the analysis provided in Section 7.4. Moreover, when the batch size exceeds $10^{3}$, the throughput of GastCoCo remains relatively stable, indicating that GastCoCo possesses excellent scalability. When the batch size is $10$ and $100$, Terrace, PCSR, and GraphOne exceed GastCoCo. This is because the construction and scheduling overheads in GastCoCo’s CBList are difficult to balance against the benefits brought by coroutines for smaller batch sizes.
Figure 14. Performance analysis. (SE stands for Sequential Execution, IE stands for Interleaved Execution, and SP stands for Software Prefetching)
@@
Figure 17. Comparison of AOA and AOE on ogbl-citation2.
Figure 18. Batch update analysis.

### v4

@@ 6.1. Execution Strategy Tuner
In this subsection, we explain how the execution strategy tuner tailors partition and scheduling strategies based on graph processing tasks and how the task allocator allocates tasks to coroutines. It is noteworthy that modeling each task as a coroutine and creating the interleaved execution mode among multiple tasks is feasible and easy. Thus, we focus on building interleaved execution patterns among multiple coroutines within a single task in this subsection. We present the logic of the execution strategy tuner in Figure 8 for a better understanding.
@@ 4 Function Scheduler($coroutine\_pool$):
9                         $coroutine\_pool[i].$destroy()
@@ 15 Function Coroutine_Func():
Graph Update Tasks. In GastCoCo, we support both individual updates and batch updates, depending on the specific requirements of the application scenario. However, it is important to note that the coroutine optimization scheme used in GastCoCo can only optimize batch edge updates. Specifically, for individual updates, we apply vertex-level locks on the CBList to ensure accuracy. For batch updates, similar to systems that support batch updates (e.g., Aspen (Dhulipala et al., 2019) and Terrace (Pandey et al., 2021)), we classify update tasks by source vertex before updating to avoid the overhead of locks caused by data conflicts. Each vertex that is waiting to be updated will have a collection of update tasks. The task allocator models each vertex’s task collection as a coroutine, which is similar to FindNeighbor($edge$) in algorithm 2. Since insertions may modify the structure making it difficult to estimate the number of suspension points before the update, we opt for the trimmed-polling scheduler for batch updates.
@@ 6.2. Hybrid Prefetching
Hardware prefetching fetches data silently to users. However, if hardware prefetching succeeds, it will lead to additional overhead to repeatedly operate software prefetching. Accordingly, the ideal scenario is to use software prefetching when the hardware prefetching fails to fetch data and to avoid using software prefetching when hardware prefetching is effective.
Assume that the cost of a cache miss is $C_{m}$ and the probability of a cache miss being resolved by hardware prefetching is $P_{h}$. This probability varies based on the memory layout of the data. The software prefetching solution via coroutines incurs a certain overhead, $C_{coro}$. If the cost $C_{m}\times(1-P_{h})\textless C_{coro}$, then the software prefetching becomes redundant. To avoid this overhead of redundant data fetching, we aim to avoid employing software prefetching on data that is highly likely to be prefetched by the hardware prefetchers. We also need to consider the phenomenon caused by the skewed graph structures from three perspectives discussed in subsection 2.2.
Therefore, we aim to use software prefetching to complement hardware prefetching for better performance. However, implementing software prefetching via coroutines also has overhead, as shown in section 5. To better utilize hardware and software prefetching, we design four prefetching strategies, as shown in Figure 9.
@@
Hybrid strategy I (Hybrid prefetching based on the block size). In the implementation of CBList, all small chunks are arranged contiguously in memory, resulting in a high probability of being prefetched by the hardware prefetcher. In this case, we skip software prefetching for small chunks and instead rely on the hardware prefetcher’s inherent capability to prefetch these small chunks. In most cases, this strategy is more effective than the previous “All Soft” strategy. This strategy is more suitable for scenarios where small chunks have a high cache hit rate.
Hybrid strategy II (Hybrid prefetching based on the hotness). To further prevent redundant data fetching, we consider the cold start issue of the jump-pointer mechanism described in Section 2.2. Based on Hybrid I, we use software prefetching at the beginning of the linked list. For the latter part of the linked list, we solely rely on the hardware prefetcher’s jump-pointer mechanism. In summary, this strategy utilizes software or hardware prefetching based on the cache’s “cold” and “hot” data status. Under ideal conditions, this strategy achieves optimal results. However, it requires iterative tuning and adjustments to determine the optimal threshold parameters of using hardware or software prefetching based on the cache’s state. Thus, the strategy is particularly suitable for scenarios where repeated fine-tuning of computations is possible.
@@ 7. EVALUATION
This section introduces the experimental setup, and then validates the efficiency of the proposed system through various evaluations.
@@ 7.1. Experimental Setup
We run the experiments on a server equipped with one Intel Xeon Platinum 8269CY CPU whose clock speed is 2.5GHz. The server contains 26 cores (52 hyper-threads) and 371GB of main memory. The sizes of the L1-cache, L2-cache, and L3-cache of the CPUs are 832KB(data cache)+832KB(instruction cache), 26MB, and 35.8MB, respectively. We compiled all systems with the O3 optimization flag using GCC v10.3. All experiments are conducted with 52 worker threads by default unless otherwise stated, with each test running five times to report the average outcome.
Datasets. We used five real-life graphs in our evaluation (see Table 2), including social networks Livejournal (Backstrom et al., 2006), Hollywood (Boldi and Vigna, 2004) and Com-friendster (Yang and Leskovec, 2012), and web graph UK-2002 (Boldi and Vigna, 2004) and Orkut (Yang and Leskovec, 2012). The detailed information of the datasets is shown in Table 2, which includes the name of the dataset, the number of vertices, the number of edges, and average degrees. To simulate the realism of dynamic graph processing scenarios in practice, the datasets are shuffled. This is because other structures that contain pointers may be penalized during loading due to the inherent order of the data. To ensure fairness, all systems load datasets using the weighted graph mode, and random weights are generated for those unweighted datasets. Besides, we used the SNB-sf1000 dataset (Angles et al., 2020) to test the system’s update performance in real-world scenarios. The SNB-sf1000 dataset includes timestamped person nodes and their relationships, along with 7,285 vertex deletions and 39,877,751 edge insertions. We also used the Ogbl-citation2 (Wang et al., 2020) dataset, which includes 128-dimensional word2vec features, to test graph attribute updates.
Workloads. We conduct three types of workloads. (i) Graph Query: we randomly query 5% edges of the data set. (ii) Graph Algorithms: we use five typical graph analysis algorithms, i.e., BFS, Single Source Shortest Path (SSSP) (Chakaravarthy et al., 2017), PageRank (PR) (Page et al., 1998), Connected Components (CC) and Label Propagation (LP). (iii) Graph Update: we synthetically generate 10,000 random graph updates in the form of edge insertions and edge deletions.
@@
LLAMA, Teseo, and PCSR utilize structures similar to CSR to enhance data locality. LiveGraph employs a log-structured approach for data insertion to achieve faster insertion performance. GraphOne and RisGraph adopt block-based adjacency lists and adjacency arrays, respectively, to support rapid data insertion. Sortledton and Terrace employ different structures for storing neighbor data based on vertex degree, striking a balance between insertion performance and graph analysis capabilities.
@@ 7.2. Performance of Graph Query
We first evaluate the performance of edge queries by GastCoCo on all datasets in Table 2, which plays a crucial role in multiple domains such as social network analysis, recommendation systems, and transportation networks. For example, in social networks, edge queries can be utilized to check whether there is a direct friendship between two users or not.
Figure 10 shows the normalized time for each system to execute queries on different datasets. Here the response time of GastCoCo is treated as the baseline, i.e., finishes in unit time. We can observe that GastCoCo outperforms others in all the cases. Specifically, GastCoCo achieves an average 13.7$\times$ (up to 16.8$\times$) speedup over Terrace, 7.4$\times$ (up to 12.0$\times$) speedup over LLAMA, 2.8$\times$ (up to 7.2$\times$) speedup over PCSR, 10.9$\times$ (up to 20.4$\times$) speedup over GraphOne, 4.7$\times$ (up to 8.5$\times$) speedup over RisGraph, 6.0$\times$ (up to 9.1$\times$) speedup over LiveGraph, 4.6$\times$ (up to 6.2$\times$) speedup over Teseo, and 5.6$\times$ (up to 7.6$\times$) speedup over Sortledton. For Terrace, LLAMA, and GraphOne, they experienced a slowdown exceeding 10$\times$ on the Hollywood dataset. This can be attributed to the dataset’s vertices having the highest average degree (i.e., 101.81), which leads to significant query overhead for vertices with high degrees due to the storage structures of these three systems during retrieval. Specifically, Terrace adopts a hierarchical structure. When the degree of a vertex is large, its neighbors will be distributed in multiple layers of data structures, which results in the need to search across multiple storage structures when querying. For GraphOne, it uses a block-based adjacency list to store the neighbors of each vertex. Therefore, for vertices with a high degree, this necessitates traversing a longer chain, thereby extending the query time. Compared with other systems, GastCoCo also achieves better performance, mainly because GastCoCo adopts more “stubby” structures, i.e., B+ Tree, as part of our edge storage. The B+ tree, due to its strict rebalance rules, has fewer levels, meaning that we encounter fewer pointers and experience fewer cache misses during queries. We can offset the benefits with software prefetching via coroutines to amortize the rebalance overhead during updates.
@@ 7.3. Performance of Graph Analysis
We next evaluate the performance of graph analysis, including the execution time of five graph analytics algorithms, by comparing GastCoCo with competitors. Figure 11 shows the normalized time for each system to execute graph analysis algorithms on different datasets. Here the response time of GastCoCo is treated as the baseline, i.e., finishes in unit time. We can see that GastCoCo outperforms others in most cases. Specifically, GastCoCo achieves an average 1.4$\times$ (up to 3.9$\times$) speedup over Terrace, 41.1$\times$ (up to 243.8$\times$) speedup over LLAMA, 11.0$\times$ (up to 39.7$\times$) speedup over PCSR, 21.7$\times$ (up to 79.4$\times$) speedup over GraphOne, 4.9$\times$ (up to 22.3$\times$) speedup over RisGraph, 4.2$\times$ (up to 16.0$\times$) speedup over LiveGraph, 5.8$\times$ (up to 21.7$\times$) speedup over Teseo, and 2.8$\times$ (up to 6.8$\times$) speedup over Sortledton. As shown in Figure 11, LLAMA and GraphOne exhibit the most significant slowdown. LLAMA employs a multi-version array to store the neighboring edges of vertices. As data is inserted, the edges of a vertex may be distributed across multiple arrays, which is disadvantageous for graph analysis algorithms since they always require access to all neighbors of a vertex. LLAMA connects the edges of a vertex distributed in different arrays by pointers, thereby reducing search time. However, this still results in extensive pointer chasing. GraphOne faces similar issues with expensive pointer chasing due to its use of block-based adjacency lists. Furthermore, Figure 11 shows that Terrace, PCSR, and Sortledton achieved better graph analysis performance compared to other competitors. This is attributed to Terrace and Sortledton using arrays to store the neighbors of vertices with low degrees, and PCSR using arrays to store the neighbors of all vertices, which have a better spatial locality. However, this design also results in higher update costs due to data movements caused by insertions/deletions, which will be detailed in Section 7.4. Additionally, Terrace and Sortledton employ B-Trees or skip lists to store the neighbors of high-degree vertices, still leading to extensive pointer chasing. In contrast, GastCoCo utilizes software prefetching and coroutines to accelerate data retrieval in graph analysis and reduce cache misses, thereby achieving superior graph analysis performance.
@@ 7.4. Performance of Graph Updates
We evaluate the graph update performance of different systems on LiveJournal, Hollywood and UK-2002 for continuous edge insertions/deletions of 10K edges. Figure 12 shows the throughput (edges per second) for all systems under different data sets. GastCoCo outperforms all competitors in all tests. Specifically, GastCoCo achieves an average 2.5$\times$ (up to 4.5$\times$) speedup over Terrace, 152.4$\times$ (up to 180.2$\times$) speedup over LLAMA, 33.3$\times$ (up to 35.4$\times$) speedup over PCSR, 4.3$\times$ (up to 5.0$\times$) speedup over GraphOne, 58.2$\times$ (up to 75.0$\times$) speedup over RisGraph, 74.4$\times$ (up to 109.4$\times$) speedup over LiveGraph, 68.8$\times$ (up to 87.2$\times$) speedup over Teseo, and 33.3$\times$ (up to 35.4$\times$) speedup over Sortledton. It can be observed that Terrace, GraphOne, and GastCoCo have higher throughput. GraphOne utilizes a block-based adjacency list storage structure, which is particularly advantageous for insertions. However, as demonstrated in the previous experiment, it performs poorly in graph queries and graph analysis. Both Terrace and GastCoCo adopt a batch updates design, i.e., classify update tasks by source vertices before updating, to avoid lock overhead caused by data conflicts. Moreover, since Terrace maintains the updated data in an ordered manner, a search operation is required to determine the position of the new data, leading to a significant number of cache misses. In contrast, although GastCoCo also maintains ordered updates, our coroutine-related design effectively reduces cache misses, thereby achieving better performance than Terrace.
@@ 7.5. Varying Batch Size of Graph Updates
We also conduct experiments to evaluate the impact of batch size of graph updates on throughput performance. Figure 13 reports the throughput of the systems on graph updates with different batch sizes on LiveJournal, Hollywood, Orkut, and UK-2002. For each batch size from $10$ to $10^{7}$, we report the throughput (edges per second) for all systems. Figure 13 shows that GastCoCo outperforms other systems with batch sizes ranging from $10^{3}$ to $10^{7}$, the reason for which is primarily consistent with the analysis provided in Section 7.4. Moreover, when the batch size exceeds $10^{3}$, the throughput of GastCoCo remains relatively stable, indicating that GastCoCo possesses excellent scalability. On the other hand, when the batch size is $10$ and $100$, Terrace, PCSR, and GraphOne exceed GastCoCo. This is because the construction and scheduling overheads in GastCoCo’s CBList are difficult to balance against the benefits brought by coroutines for smaller batch sizes.
Figure 14. Performance analysis. (SE represents Sequential Execution, IE represents Interleaved Execution, and SP represents Software Prefetching)